OpenStudy (anonymous):

Help with horizontal distance, initial velocity and angles problem? Medal and fan 17. Recall that the equation for the horizontal distance h in feet of a projectile with initial velocity v0 and initial angle theta is given by h=v0^2/16 sin theta cos theta. a. Assume the initial velocity is 60 feet/second . What initial angle will you need to ensure that the horizontal distance will be exactly 100 feet? (3 points) b. Assume the initial velocity is 60 feet/second. What is the maximum horizontal distance possible and at what angle does this occur?

OpenStudy (anonymous):

Could you help me with this one? @zepdrix

zepdrix (zepdrix):

I'm not familiar with that equation. Are the sine and cosine in the denominator or no? :d

OpenStudy (anonymous):

No, they aren't. They're next to it.

zepdrix (zepdrix):

\[\Large\rm h=\frac{v_{o}^2}{16}\sin \theta \cos \theta\]k

zepdrix (zepdrix):

For part a, we let \(\Large\rm v_{o}=60\) and \(\Large\rm h=100\) and try to solve for our angle \(\Large\rm \theta\), yes?

OpenStudy (anonymous):

Yes. 100 = (60)^2/ 16 sin theta cos theta

zepdrix (zepdrix):

Remember your Sine Double Angle Identity? :)

OpenStudy (anonymous):

sin(2x) = 2sinxcosx?

zepdrix (zepdrix):

\[\Large\rm 100=\frac{(60)^2}{16}\color{orangered}{\sin \theta \cos \theta}\]Ok good. So we need a 2 right? Hmmm where can we get a 2 from?

OpenStudy (anonymous):

Just multiply both sides by 2?

zepdrix (zepdrix):

Yah there we go :3\[\Large\rm 200=\frac{(60)^2}{16}\color{orangered}{2\sin \theta \cos \theta}\]

zepdrix (zepdrix):

\[\Large\rm 200=\frac{(60)^2}{16}\color{orangered}{\sin(2\theta)}\]

OpenStudy (anonymous):

Wouldn't you multiply the (60)^2/16 as well?

zepdrix (zepdrix):

Multiply that to the other side? Yah seems like a good idea.

zepdrix (zepdrix):

\[\Large\rm \frac{200\cdot 16}{60^2}=\color{orangered}{\sin(2\theta)}\]

zepdrix (zepdrix):

Ah sorry were you typing it? >.< Did I steal the fun? lol

zepdrix (zepdrix):

I gotta learn to slow down sometimes :P heh

OpenStudy (anonymous):

Noo, I meant multiplying it by the 2.

zepdrix (zepdrix):

Oh oh oh. No, we only multiply the right side by 2 `one time`. Recall multiplication:\[\Large\rm 2(3x)\quad \ne\quad 2\cdot3\cdot 2\cdot x\]We don't give a 2 to each in this example.

zepdrix (zepdrix):

Don't confuse multiplication with exponents! We certainly do ... do that with exponents.\[\Large\rm (3x)^2\quad=\quad 3^2\cdot x^2\]

zepdrix (zepdrix):

I gave a 2 to the right side already. See how I put it next to the sin theta?

zepdrix (zepdrix):

Or did you mean something else +_+

OpenStudy (anonymous):

Noo, you're right. Sorry I'm tired.

OpenStudy (anonymous):

Okay, I gotcha so far.

zepdrix (zepdrix):

So far we have:\[\Large\rm number=\sin(2\theta)\]

zepdrix (zepdrix):

Understand how we can solve for theta from here? Remember how to use your inverse sine function and stuff?

OpenStudy (anonymous):

Sorry, I'm trying to remember currently. That was a bit ago.

OpenStudy (anonymous):

Wouldn't it be x= 1/2 sin^-1y?

zepdrix (zepdrix):

\[\Large\rm stuff=\sin(\text{angle)}\]Remember with inverse functions we swap the x and y? Inverse trig functions are the same idea, we swap the stuff with the angle.\[\Large\rm \text{angle}=\sin^{-1}(stuff)\]

OpenStudy (anonymous):

Okay...was my answer right?

zepdrix (zepdrix):

I dunno I'm not sure what you were doing >.< Where is the 1/2 coming from? Is that from the 2 that was connected to theta?

OpenStudy (anonymous):

Yeah, it was. I'm sorry, I can't remember that at all for some reason. I don't understand how you can have arcsine 2 theta when it's not even in the range..

zepdrix (zepdrix):

No that's fine :) I guess you were just moving ahead. I got confused because you changed the letters to x and y.\[\Large\rm number=\sin(2\theta)\]Applying the inverse,\[\Large\rm 2\theta=\sin^{-1}(number)\]Dividing by 2,\[\Large\rm \theta=\frac{1}{2}\sin^{-1}(number)\]

zepdrix (zepdrix):

I got a little lazy with our big messy number... let's see it was ummm...\[\Large\rm number=\frac{200\cdot16}{60^2}\]

zepdrix (zepdrix):

So we need a calculator at this point :o

OpenStudy (anonymous):

Yeahhh, that's where I was going, sorry!

OpenStudy (anonymous):

Okay, hold on. Are we multiplying 1/2 arcsin by 2 then by sin again?

zepdrix (zepdrix):

Whut? :U No. We're not involving sine anymore. Only arcsine.

OpenStudy (anonymous):

So angle 30...are we plugging that back into the original sinthetacostheta?

zepdrix (zepdrix):

That answers part a.

OpenStudy (anonymous):

So we're working on b now? The answer to A was theta = 30?

zepdrix (zepdrix):

Umm I came up with 31.4 degrees.

OpenStudy (anonymous):

Wait for what?? Cause arcsin 1.2 is 30 degrees.

OpenStudy (anonymous):

1/2*

zepdrix (zepdrix):

\(\Large\rm \dfrac{200\cdot16}{60^2}\approx 0.888889\) \[\Large\rm \frac{1}{2}\arcsin(0.888889)\approx 31.36697778\]

zepdrix (zepdrix):

The 1/2 is OUTSIDE of the arcsine. :o we're taking the arcsine of our big messy number.

zepdrix (zepdrix):

And then multiplying that result by 1/2

zepdrix (zepdrix):

Get confused in there somewhere? :o

OpenStudy (anonymous):

Yeah haha, now I understand why we used the inverse instead of regular sin. Okay, so the number that is achieved for theta is 31.4?

zepdrix (zepdrix):

Mmmm yah looks that way >.<

OpenStudy (anonymous):

Alright awesome. So for b, we don't have h=100 and our angles aren't given again..

zepdrix (zepdrix):

Part b is a little trickier.\[\Large\rm h=\frac{60^2}{16}\sin \theta \cos \theta\]We want to maximize our h.

OpenStudy (anonymous):

How would we maximize it?

zepdrix (zepdrix):

I'm gonna do something a little sneaky so we can write this as a single trig function again. Instead of multiplying each side by 2, I'm going to multiply the right side by 2 and 1/2 ok?\[\Large\rm h=\frac{1}{2}\cdot\frac{60^2}{16}\cdot2\sin \theta \cos \theta\]

zepdrix (zepdrix):

We need that 2 for our Sine Double Angle Identity. The 2 in the denominator we'll just combine with the 16.\[\Large\rm h=\frac{60^2}{32}\color{orangered}{2\sin \theta \cos \theta}\]

zepdrix (zepdrix):

Was that step ok? I was essentially giving the right side a 2/2 and placing each 2 in a different spot.

OpenStudy (anonymous):

Yeah that makes sense!

zepdrix (zepdrix):

\[\Large\rm h=\frac{60^2}{32}\color{orangered}{\sin2\theta}\]

OpenStudy (anonymous):

Right on, so we'll use the same 1/2 arcsin?

zepdrix (zepdrix):

No no no. We would use the inverse function if we needed an `angle`. Here we need height. So we need to do something different.

zepdrix (zepdrix):

Do you remember what the `amplitude` of a sine function tells us?\[\Large\rm y=A \sin x\] \(\Large\rm A=amplitude\)

OpenStudy (anonymous):

Yeah..would that mean that the big number is the amplitude?

zepdrix (zepdrix):

The amplitude tells us how `big` our function can get: and max and min values. Yes, that coefficient in front of our sine is our amplitude.

zepdrix (zepdrix):

That number will be our maximum height. We need to figure out `where` it happens though.

zepdrix (zepdrix):

\(\Large\rm h=112.5\color{orangered}{\sin2\theta}\) So we have something like this, yes? And we want to know what `angle` gives us our maximum, 112.5. In other words, let h=112.5 and solve for our angle again.

OpenStudy (anonymous):

Ohhhh, alright. So would we plug in 112.5 into 1/2 arcsin?

zepdrix (zepdrix):

No. You're plugging 112.5 into h.

zepdrix (zepdrix):

\[\Large\rm 112.5=112.5\sin 2 \theta\]

OpenStudy (anonymous):

Ohhhhhhhh, then you would divide and the left side would be 1?

zepdrix (zepdrix):

Mmm good!\[\Large\rm 1=\sin 2\theta\]

OpenStudy (anonymous):

So the angle is 45?

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

Awesome, thanks so much(: Sorry I was so slow, I'm working on my AP Chem assignment at the same time. I somehow turned out to excel at chemistry but definitely not math lol

zepdrix (zepdrix):

hehe XD